OpenStudy (anonymous):

√2x = -5 i got no solution am i right? √5x + 3 =8 solving for x I get 61/5 so is the answer 5?

OpenStudy (anonymous):

i get sqrt(2) sqrt(x) = -5

OpenStudy (anonymous):

isn't it 25/2 ?

OpenStudy (p0sitr0n):

5x^1/2 = 5 5x = 25 x = 5

OpenStudy (anonymous):

might be

OpenStudy (anonymous):

is it root(2)x or root(2x) ?

OpenStudy (anonymous):

(2x)

OpenStudy (anonymous):

the second one is 5

OpenStudy (anonymous):

pretty sure the first one is 25/2

OpenStudy (anonymous):

root(2x) = -5 square both sides =2x = 25 divide both sides by 2 x = 25/2

OpenStudy (anonymous):

root(5x) + 3 = 8 is 5

OpenStudy (anonymous):

yeah ur right

OpenStudy (anonymous):

subtract 3 both sides root(5x) = 5 square both sides 5x = 25 divide by 5 x = 5

OpenStudy (anonymous):

yeah got 25/2 thank you

OpenStudy (anonymous):

yw